You have taken 3 tests and have an average of 72 points. If you score 100 points on each of the rest of your tests, how many more tests do you need to take to raise your average to 88 points?
Answer by checkley77(12569)

(3*72+x*100)/(x+3)=88
(216+100x)/(x+3)=88
216+100x=88(x+3)
216+100x=88x+264
100x-88x=264-216
12x=48
x=48/12
x=4 more tests with grades of 100 should do it.
Proof:
(216+4*100)/(4+3)=88
(216+400)/7=88
616/7=88
88=88